# Causal Climate Emulation with Bayesian Filtering

### v1

@@ 4.3 Learning a causal model improves PICABU’s generalization performance
On pre-industrial data, we train PICABU and two ablated models, where the sparsity and orthogonality constraints are separately dropped from the loss (Appendix L). We evaluate these models by looking at their next timestep predictions of surface temperature when provided with 256 random initial conditions from three different emissions scenarios, sampled from SSP2-4.5, SSP3-7.0, and SSP5-8.5 between years 2070 and 2100. These correspond to a considerably warmer climate than the pre-industrial data used for training. Note that we do not include the anthropogenic emissions as inputs to the emulator; instead, these experiments evaluate the ability of the models to generalize to a new distribution of initial conditions. Table 3 shows the mean absolute error (MAE), coefficient of determination (R2), and LSD for next timestep predictions.
@@
No sparsity $**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$$**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$$**{}^{\textbf{*}}$ 0.048 $**{}^{\textbf{*}}$$**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$
No ortho $**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$$**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$$**{}^{\textbf{*}}$ 0.049 $**{}^{\textbf{*}}$$**{}^{\textbf{*}}$ $**{}^{\textbf{*}}$
Table 3: Learning a causal graph helps generalization to new emission scenarios. Next-timestep prediction results when given initial conditions from out-of-distribution data. For each scenario, 3 metrics are shown: MAE (lower is better), $R^{2}$ (higher is better), and LSD (lower is better). We evaluate PICABU, PICABU without the sparsity constraint, and PICABU without the orthogonality constraint with initial conditions corresponding to scenarios SSP2-4.5, SSP3-7.0, and SSP5-8.5 (from left to right, diverging further from the training distribution). indicates a statistically significant difference, obtained using a t-test. For $R^{2}$, the relative difference is taken with respect to 1, reflecting how much further the metric is from the ideal $R^{2}=1$.
@@ 4.4 Counterfactual experiments for attribution of extreme events
Climate models are crucial for attribution studies. When large anomalies occur, e.g. a particularly hot year, climate scientists typically perturb the inputs to ESMs and observe any changes in the projected climate to attribute causal drivers to these events (Kosaka and Xie, 2013). This is compute- and time-consuming, particularly if many experiments are needed to capture the probabilistic nature of the simulations. With PICABU, due to its causal interpretability, we can directly intervene on our learned causal graph to carry out counterfactual experiments.
As a case study, we consider the anomalously warm series of months around November 1851 in the NorESM2 data. We identify the latent variable in our model that controls sea surface temperatures in the equatorial east Pacific, a widely used indicator of ENSO. This is possible because the model learns latent variables that are mapped to distinct regions of the world following the single-parent assumption (Figure 4). We then carry out a counterfactual experiment: we intervene on this variable and manually set it to a new, higher value, and observe the resulting change in the next timestep prediction for November 1851. No other alteration to the latents or causal graph is made.
This intervened variable influences other latents through the learned causal graph, which is used to generate next timestep latents, which are in turn decoded to produce temperature fields. Figure 4 shows that if El Niño had been stronger, then temperatures would have been even higher globally, according to the learned causal model. Unlike traditional models, generating multiple counterfactual projections is simple and rapid in our framework.
Figure 4: Intervention on the latent variable that controls the ENSO state. The left panel shows mapping from latents to observations (colored grid points), and the intervened latent. We increase its value, which influences other latents through the learned causal graph. No other variables are intervened on. The middle panel shows the original, unintervened next-step prediction, and the right panel shows the intervened prediction.
@@ 5 Discussion
Limitations and future work. The main limitation of our model arises from the single-parent decoding assumption, which effectively partitions the observation space into subspaces mapped to each latent. A high latent dimension is required to represent certain complex prediction spaces, which may contribute to the high-frequency noise we observe in GMST. While single-parent decoding remains a necessary assumption to enable the recovery of causal structure under our framework, relaxing this assumption would be a promising area of future inquiry. On the climate science side, considering anthropogenic emissions, and allowing the model to represent connections between the different variables in the climate system, are natural next steps for improved emulation of climate model data under different emission scenarios. Training the model to attribute present-day extreme events to climate change (using reanalysis data) is also a promising avenue for future work.
@@ 6 Acknowledgements
This project was supported by the Intel-Mila partnership program, IVADO, Canada CIFAR AI Chairs program, and Canada First Research Excellence Fund. We acknowledge computational support from Mila – Quebec AI Institute, including in-kind support from Nvidia Corporation. P.N. was supported by the UK Natural Environment ResearchCouncil (Grant NE/V012045/1). A.T.A. was financially supported by NERC through NCAS (R8/H12/83/003). The authors thank Charlotte E.E. Lange for helping to build the competing methods.
@@ Appendix A Data availability
Processed data used in this work are available at an anonymous Zenodo repository (https://zenodo.org/records/14773929). Raw NorESM2 data can be downloaded from the ESGF CMIP6 data store.
@@ B.1 CDSD generative model
The model considers a stationary time series of order $\tau$ over these latent variables. Binary matrices $\left\{G^{k}\right\}_{k=1}^{\tau}$ represent the causal relationships between latents at different time steps. Specifically, an element $G^{k}_{ij}=1$ indicates that $z^{t-k}_{j}$ is a causal parent of $z^{t}_{i}$, capturing the lagged causal relations between the time-steps $t-k$ and $t$. The adjacency matrix $F$ delineates the causal connections between the latent variables $\mathbf{z}$ and the observed variables $\mathbf{x}$. Each observed variable $x_{i}$ has at most one latent parent, adhering to the single-parent decoding structure.
@@
[…]
@@
The graph between the latent $\mathbf{z}$ and the observable $\mathbf{x}$ is parameterized using a weighted adjacency matrix $W$. To enforce the single-parent decoding, $W$ is constrained to be non-negative with orthonormal columns. Neural networks are optionally used to parameterize encoding and decoding functions $g_{j}$, $f_{j}$, $\tilde{\mathbf{f}}$. The graphs $G^{k}$ are sampled from $G^{k}_{ij}\sim Bernoulli(\sigma(\Gamma^{k}_{ij}))$, with $\Gamma^{k}$ being learnable parameters. The objective is optimized using stochastic gradient descent, leveraging the Straight-Through Gumbel estimator  [Jang et al., 2017] and the reparameterization trick [Kingma and Welling, 2014].
@@ B.2 Inference with CDSD: Objective and Optimization
In this section, we present how inference and optimization is carried out when using CDSD [Brouillard et al., 2024].
@@
$\max_{W,\Gamma,\mathbf{\phi}}\mathbb{E}_{G\sim\sigma(\Gamma)}\bigl{[}\mathbb{E}_{\mathbf{x}}\left[\mathcal{L}_{\mathbf{x}}(W,\Gamma,\mathbf{\phi})\right]\bigr{]}-\lambda_{s}||\sigma(\Gamma)||_{1}$ (12)
@@
$\mathcal{L}_{\mathbf{x}}$ is the ELBO corresponding to the right-hand side term in Equation 11 and $\lambda_{s}>0$ a coefficient for the regularization of the graph sparsity. The non-negativity of $W$ is enforced using the projected gradient on $\mathbb{R}_{\geq 0}$, and its orthogonality enforced using the following constraint:
@@
$\max_{W,\Gamma,\mathbf{\phi}}\mathbb{E}_{G\sim\sigma(\Gamma)}\bigl{[}\mathbb{E}_{\mathbf{x}}[\mathcal{L}_{\mathbf{x}}(W,\Gamma,\mathbf{\phi})]\bigr{]}\\ -\lambda_{s}||\sigma(\Gamma)||_{1}-\text{Tr}\left(\lambda_{W}^{T}h(W)\right)-\frac{\mu_{W}}{2}||h(W)||^{2}_{2},$ (13)
@@ B.3 Parameter sharing
The original model parameterizes the distribution of the observations with one neural network per grid point and variables, and the transition model with one neural network per latent variable, leading to the number of neural networks linearly increasing with the number of latents and input dimensions. We implement parameter sharing for computational efficiency. More precisely, there is now a single neural network to parameterize the mapping and a single neural network to parameterize the transition model. The neural networks take as input a positional embedding that is learned through training, and represents each grid location and input variable (non-linear mapping from latents to observations), and each latent (non-linear transition model).
@@ Appendix C SAVAR datasets
C.1 Details on the training procedure for all models
@@
PICABU was trained using the parameters shown in Table 6. The only difference was the sparsity constraint “Constrained value", which was set to match the expectation of the causal graph, as explained in Section 4.1.
CDSD was trained using the same parameters, except that instead of a sparsity constraint, it uses a penalty. The coefficient $\mu$ was thus fixed. We performed a hyperparameter search over $\mu$ and tested values $10^{-16},10^{-14},...,10^{2}$. We retained the value leading to the highest F1 score. Moreover, we stopped the training as the sparsity was reaching its expected value.
PCMCI was trained using ParCorr to perform the conditional independence tests, and the significance value of the tests was replaced to keep the $M$ most significant instead, with $M$ being the expected value of edges in the graph.
An advantage of PICABU is the possibility to control the final sparsity of the graph directly, as it is optimized using a constraint. We thus set the sparsity equal to the expected number of edges in the true graph $M$. In comparison, CDSD utilizes a penalty to optimize for sparsity. One drawback is that if the penalty coefficient is too high or training lasts too long, the learned graph will be too sparse. For fair comparison, we stop the training of CDSD when the number of connections of strength at least $0.5$ is equal to $M$. We train with various values for the sparsity penalty coefficient, and report the maximum F1 score obtained. PCMCI instead computes conditional independence tests between all possible pairs of latent variables, and keeps edges when the test rejects conditional independence. The test is parameterized by a threshold for p-values: if the test’s p-value is above the threshold, conditional independence is rejected. For fair comparison, we modify its procedure to keep links for the $M$ conditional independence tests with the highest p-values. The final parameters used for the three methods are given in Appendix C.
C.2 Inferring causally-relevant latents improves accuracy
Figure 7: PICABU disentangles latents well. The top row shows the four modes of variability identified by PICABU, in the dataset shown in Figure 6 (A). The bottom row shows the four modes identified by Varimax-PCMCI. The latents are not very well disentangled, as they are inferred independently of the causal graph, and PCA exhibits timescale mixing and exhibit timescale mixing Aubry et al. [1993]. Varimax-PCA is not able to disentangle them because of the autoregressive dependencies present in the data.
@@ Appendix D Training details for ViT + positional encoding model
The ViT + pos. encoding model is adapted from ClimaX [Nguyen et al., 2023]. This architecture is originally designed to take multiple variables as input. We adapted it to only emulate temperature. We trained it with various hyperparameters and report results with the best test loss. We tried different network depths (from 1 to 8) and got best results with 2, various learning rates (from 0.00001 to 0.1) and got best results with 0.0001, various batch sizes (from 16 to 512) and got best results with 64. The best architecture is significantly smaller than the original one since it is trained to emulate one variable. Moreover, it is trained to predict multiple timesteps into the future, by having a fixed encoder and multiple prediction heads, each corresponding to a different prediction horizon. Here, we train it to predict 1, 2, 3, 4, or 5 timesteps into the future to compare to PICABU with $\tau=5$. During rollouts, we select the prediction head corresponding to $\tau=1$ to emulate monthly trajectories.
The architecture is adapted from an approach designed to be trained with more data sources and input variables, and at multiple time resolution. We find that when it is trained on a smaller dataset, it is unable to learn the dynamics of the system directly. The architecture is trained with a latitude-weighted mean squared error [Rasp et al., 2020], which will suffer from learning low spatial frequencies. Hence, the model has low bias but very high range and variability.
@@ Appendix E Example time series of NorESM2 data
Figure 8 shows an example time series of NorESM2 data, illustrating six consecutive months of data. These data have been normalized and deseasonalized. This time series illustrates important features of seasonal and climate variability. For example, while the Pacific Ocean shows relatively unchanging temperatures, temperatures in the extra-tropics are much more variable and less predictable. Note that for our experiments on NorESM2 data, PICABU takes as input the previous five months (here, $t-5$ to $t-1$), and then learns to predict timestep $t$ based on these past timesteps.
@@ Appendix F Further climate emulation results
To describe the performance of PICABU for long rollouts, we illustrate an example time series of the Niño3.4 index over a 50-year rollout, and similarly, we show the time series of normalized global mean surface temperature in Figure 9. These rollouts are generated from a randomly chosen initial condition. The ENSO variability generated by the emulator is realistic, capturing the distribution of observed variability well without drift. While the GMST similarly does not drift away from the expected mean value, our emulator exhibits too much variance over short timescales.
@@
This observation is confirmed by the power spectra for ENSO and GMST. While the ENSO power spectrum generated by PICABU is accurate at both low and high frequencies, the power for GMST is too high at high frequencies, as shown in Figure 10. This is true across all variants of PICABU that we train. The V-PCMCI model, however, has a more accurate spectrum at high frequencies but very little power at low frequencies. We also show the spectra at higher frequencies, on a log scale in Figure 11, where it is clear that PICABU captures the high frequencies of ENSO much better than for GMST.
Figure 10: PICABU learns accurate temporal variability for ENSO, and outperforms other methods in learning GMST variability. We run PICABU for ten different 50-year emulations, and compute the mean and standard deviation of the spectra, doing the same for ten different 50-year periods of NorESM2 data. On the left, the power spectra for the Niño3.4 index, for PICABU and ablations, and ground-truth data (NorESM2) is shown. On the right, the same is shown for GMST.

### v2

@@ 4.3 Learning a causal model improves PICABU’s generalization performance
On pre-industrial data, we train PICABU and two ablated models, where the sparsity and orthogonality constraints are separately dropped from the loss (Appendix˜M). We evaluate these models by examining their next-timestep predictions of surface temperature when provided with 256 random initial conditions from three different emissions scenarios, sampled from SSP2-4.5, SSP3-7.0, and SSP5-8.5 between years 2070 and 2100. These correspond to a considerably warmer climate than the pre-industrial data used for training. Note that we do not include the anthropogenic emissions as inputs to the emulator; instead, these experiments evaluate the ability of the models to generalize to a new distribution of initial conditions. Table˜3 shows the mean absolute error (MAE), coefficient of determination (R2), and LSD for next timestep predictions.
@@
[…]
Table 3: Learning a causal graph helps generalization to new emission scenarios. Next-timestep prediction results when given initial conditions from out-of-distribution data. For each scenario, three metrics are shown: MAE (lower is better), $R^{2}$ (higher is better), and LSD (lower is better). We evaluate PICABU, PICABU without the sparsity constraint, and PICABU without the orthogonality constraint with initial conditions corresponding to scenarios SSP2-4.5, SSP3-7.0, and SSP5-8.5 (from left to right, diverging further from the training distribution). indicates a statistically significant difference with the corresponding PICABU score, obtained using a t-test. For $R^{2}$, the relative difference is taken with respect to 1, reflecting how much further the metric is from the ideal $R^{2}=1$.
@@ 4.4 Counterfactual experiments for attribution of extreme events
Climate models are crucial for attribution studies. When large anomalies occur, e.g. a particularly hot year, climate scientists typically perturb the inputs to ESMs and observe any changes in the projected climate to attribute causal drivers to these events (Kosaka and Xie, 2013). This is compute- and time-consuming, particularly if many experiments are needed to capture the probabilistic nature of the simulations. With PICABU, due to its causal interpretability, we can directly intervene on the input data to carry out counterfactual experiments. As a case study, we consider the anomalously warm series of months around November 1851 in the NorESM2 data. We identify the Niño 3.4 index grid cells that correspond to sea surface temperatures in the equatorial eastern Pacific, a widely used indicator of ENSO. We then carry out a counterfactual experiment: we intervene on these grid cells, manually setting their values to new, perturbed values, and observe the resulting change in the next-timestep prediction for November 1851. No other alteration to the observational space, latent variables or causal graph is made.
These intervened grid cells influence the value of their parent latent variables, which then interact through the learned causal graph to generate latents at the next timestep, which are in turn decoded to produce temperature fields. Figure˜5 shows that if El Niño had been stronger, then temperatures would have been even higher globally, according to the learned causal model. In Appendix˜P, we illustrate additional interventions on Alaskan and IOD temperatures. We also intervene in latent space rather than the observation space, which is possible as latents map to distinct regions of the world following the single-parent assumption. We show interventions on the latent variable that controls equatorial Pacific temperatures, and the direct correspondence between latent-space values and observation-space temperatures (Figure˜26). Unlike traditional models, generating multiple counterfactual projections is simple and rapid in our framework.
Figure 5: Intervention on the grid cells that describe the ENSO state. The left panel shows mapping from latents to observations (colored grid points), and the intervened grid cells. We increase their value, which then influences the latents at the current and next timestep through the learned causal graph. No other grid cells or latent variables are intervened on. The middle panel shows the original, unintervened next-step prediction, and the right panel shows the intervened prediction.
@@ 5 Discussion
Limitations and future work. The main limitation of our model arises from the single-parent decoding assumption, which effectively partitions the observation space into subspaces mapped to each latent. A high latent dimension is required to represent certain complex prediction spaces, which may contribute to the high-frequency noise we observe in GMST. While single-parent decoding remains a necessary assumption to enable the recovery of causal structure under our framework, relaxing this assumption would be a promising area of future inquiry. On the climate science side, considering anthropogenic emissions and allowing the model to represent connections between the different variables in the climate system are natural next steps for improved emulation of climate model data under different emission scenarios. Training the model to attribute present-day extreme events to climate change (using reanalysis data) is also a promising avenue for future work.
@@ 6 Acknowledgements
This project was supported by the Intel-Mila partnership program, IVADO, Canada CIFAR AI Chairs program, and Canada First Research Excellence Fund. We acknowledge computational support from Mila – Quebec AI Institute, including in-kind support from Nvidia Corporation. S.H. was supported by funding from EPSRC (EP/S022961/1). P.N. was supported by the UK Natural Environment Research Council (Grant NE/V012045/1). A.T.A. was financially supported by NERC through NCAS (R8/H12/83/003). The authors thank Charlotte E.E. Lange for helping to build the competing methods.
@@ Appendix A Data availability
Processed data used in this work are available at a Zenodo repository (https://zenodo.org/records/14773929). Raw NorESM2 data can be downloaded from the ESGF CMIP6 data store.
@@ B.1 CDSD generative model
The model considers a stationary time series of order $\tau$ over these latent variables. Binary matrices $\left\{G^{k}\right\}_{k=1}^{\tau}$ represent the causal relationships between latents at different time steps. Specifically, an element $G^{k}_{ij}=1$ indicates that $z^{t-k}_{j}$ is a causal parent of $z^{t}_{i}$, capturing the lagged causal relations between the time-steps tkt-k and $t$. The adjacency matrix $F$ delineates the causal connections between the latent variables $\mathbf{z}$ and the observed variables $\mathbf{x}$. Each observed variable $x_{i}$ has at most one latent parent, adhering to the single-parent decoding structure.
@@
Maximizing $p(\mathbf{x}^{\leq T})=\int p(\mathbf{x}^{\leq T},\mathbf{z}^{\leq T})\,d\mathbf{z}^{\leq T}$ unfortunately involves an intractable integral, hence the model is fit by maximizing an evidence lower bound (ELBO) [Kingma and Welling, 2014, Girin et al., 2021] for $p(\mathbf{x}^{\leq T})$. The variational approximation of the posterior $p(\mathbf{z}^{\leq T}\mid\mathbf{x}^{\leq T})$ is $q(\mathbf{z}^{\leq T}\mid\mathbf{x}^{\leq T})$.
@@
The graph between the latent $\mathbf{z}$ and the observable $\mathbf{x}$ is parameterized using a weighted adjacency matrix $W$. To enforce the single-parent decoding, $W$ is constrained to be non-negative with orthonormal columns. Neural networks are optionally used to parameterize encoding and decoding functions $g_{j}$, $_{j}, 𝐟~\tilde{\mathbf{f}}$. The graphs $G^{k}$ are sampled from $G^{k}_{ij}\sim Bernoulli(\sigma(\Gamma^{k}_{ij}))$, with $\Gamma^{k}$ being learnable parameters. The objective is optimized using stochastic gradient descent, leveraging the Straight-Through Gumbel estimator  [Jang et al., 2017] and the reparameterization trick [Kingma and Welling, 2014].
@@ B.2 Inference with CDSD: Objective and Optimization
In this section, we present how inference and optimization are carried out when using CDSD [Brouillard et al., 2024].
@@
$\text{s.t.}\ \ W\ \text{is orthogonal and non-negative},$
$\mathcal{L}_{\mathbf{x}}$ is the ELBO corresponding to the right-hand side term in Equation˜11 and $\lambda_{s}>0$ a coefficient for the regularization of the graph sparsity. The non-negativity of $W$ is enforced using the projected gradient on $\mathbb{R}_{\geq 0}$, and its orthogonality enforced using the following constraint:
@@
[…]
@@ B.3 Parameter sharing
The original model parameterizes the distribution of the observations with one neural network per grid point and variables, and the transition model with one neural network per latent variable, leading to a linear increase in neural networks with the number of latents and input dimensions. We implement parameter sharing for computational efficiency. More precisely, there is now a single neural network to parameterize the mapping and a single neural network to parameterize the transition model. The neural networks take as input a positional embedding that is learned through training, and represents each grid location and input variable (non-linear mapping from latents to observations), and each latent (non-linear transition model).
@@ Appendix C SAVAR datasets
C.1 Distribution of links strengths
Figure 8: Probability distribution function of the distribution of SAVAR link strengths. The beta distribution $\mathcal{B}(4,8)$ was chosen to have mostly moderate link strengths, centered around $1/3$, with few values very close to 0 and 1 and relatively low standard deviation ($0.13$).
C.2 Details on the training procedure for all models
@@
PICABU was trained using the parameters shown in Table˜7. The only difference was the sparsity constraint “Constrained value", which was set to match the expectation of the causal graph, as explained in Section˜4.1.
CDSD was trained using the same parameters, except that instead of a sparsity constraint, it uses a penalty. The coefficient $\mu$ was thus fixed. We performed a hyperparameter search over $\mu$ and tested values $10^{-16},10^{-14},...,10^{2}$. We retained the value leading to the highest F1 score. Moreover, we stopped the training when the sparsity was reaching its expected value.
PCMCI was trained using ParCorr to perform the conditional independence tests, and the significance value of the tests was replaced to keep the $M$ most significant instead, where $M$ is the expected number of edges in the graph.
An advantage of PICABU is the possibility to control the final sparsity of the graph directly, as it is optimized using a constraint. We thus set the sparsity equal to the expected number of edges in the true graph $M$. In comparison, CDSD utilizes a penalty to optimize for sparsity. One drawback is that if the penalty coefficient is too high or training lasts too long, the learned graph will be too sparse. For fair comparison, we stop the training of CDSD when the number of connections of strength at least $0.5$ is equal to $M$. We train with various values for the sparsity penalty coefficient, and report the maximum F1 score obtained. PCMCI instead computes conditional independence tests between all possible pairs of latent variables, and keeps edges when the test rejects conditional independence. The test is parameterized by a threshold for p-values: if the test’s p-value is above the threshold, conditional independence is rejected. For fair comparison, we modify its procedure to keep links for the $M$ conditional independence tests with the highest p-values. The final parameters used for the three methods are given in Appendix˜C.
C.3 Inferring causally-relevant latents improves accuracy
Figure 9: PICABU disentangles latents well. The top row shows the four modes of variability identified by PICABU, in the dataset shown in Figure˜7 (A). The bottom row shows the four modes identified by Varimax-PCMCI. The latents are not very well disentangled, as they are inferred independently of the causal graph, and PCA exhibits timescale mixing and exhibit timescale mixing Aubry et al. [1993]. Varimax-PCA is not able to disentangle them because of the autoregressive dependencies present in the data.
@@ Appendix D Training details for ViT + positional encoding model
The ViT + pos. encoding model is adapted from ClimaX [Nguyen et al., 2023]. This architecture is originally designed to take multiple variables as input. We adapted it to only emulate temperature. We trained it with various hyperparameters and report results with the best test loss. We tried different network depths (from 1 to 8) and got best results with 2, various learning rates (from 0.00001 to 0.1) and got best results with 0.0001, various batch sizes (from 16 to 512) and got best results with 64. The best architecture is significantly smaller than the original one since it is trained to emulate one variable. Moreover, it is trained to predict multiple timesteps into the future by having a fixed encoder and multiple prediction heads, each corresponding to a different prediction horizon. Here, we train it to predict 1, 2, 3, 4, or 5 timesteps into the future to compare to PICABU with $\tau=5$. During rollouts, we select the prediction head corresponding to $\tau=1$ to emulate monthly trajectories.
The architecture is adapted from an approach designed to be trained with more data sources and input variables, and at multiple time resolutions. We find that when trained on a smaller dataset, it is unable to learn the dynamics of the system directly. The architecture is trained with a latitude-weighted mean squared error [Rasp et al., 2020], which will suffer from learning low spatial frequencies. Hence, the model has low bias but very high range and variability.
@@ Appendix E Example time series of NorESM2 data
Figure˜10 shows an example time series of NorESM2 data, illustrating six consecutive months of data. These data have been normalized and deseasonalized. This time series illustrates important features of seasonal and climate variability. For example, while the Pacific Ocean shows relatively unchanging temperatures, temperatures in the extratropics are much more variable and less predictable. Note that for our experiments on NorESM2 data, PICABU takes as input the previous five months (here, t5t-5 to t1t-1), and then learns to predict timestep $t$ based on these past timesteps.
@@ Appendix F Further climate emulation results
To describe the performance of PICABU for long rollouts, we illustrate an example time series of the Niño3.4 index over a 50-year rollout, and similarly, we show the time series of normalized global mean surface temperature in Figure˜11. These rollouts are generated from a randomly chosen initial condition. The ENSO variability generated by the emulator is realistic, capturing the distribution of observed variability well without drift. While the GMST similarly does not drift away from the expected mean value, our emulator exhibits too much variance over short timescales.
@@
This observation is confirmed by the power spectra for ENSO and GMST. While the ENSO power spectrum generated by PICABU is accurate at both low and high frequencies, the power for GMST is too high at high frequencies, as shown in Figure˜4. This is true across all variants of PICABU that we train. The V-PCMCI model, however, has a more accurate spectrum at high frequencies but very little power at low frequencies. We also show the spectra at higher frequencies, on a log scale in Figure˜12, where it is clear that PICABU captures the high frequencies of ENSO much better than for GMST.